OpenStudy (anonymous):

h=2f+3g when f(x)=x^2+2x-2, g(x)=x^2+2x-2

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

\[2(x^2+2x-2)+3(x^2+2x-2)\] is a start then a raft of algebra but it looks like a strange question, because in this case \(f=g\)

OpenStudy (anonymous):

you have \[f(x)=x^2+2x-2\] and also \[g(x)=x^2+2x-2\] so why don't they call them both \(f\)?

OpenStudy (anonymous):

let me send you the document @satellite73

OpenStudy (anonymous):